OpenStudy (anonymous):

What is the solution to the equation? 1/10 + 1/5=

OpenStudy (cwrw238):

first change the 1/5 to tenths then add the numerators

OpenStudy (anonymous):

1/10+1/5 =1/10+2/10 =3/10